OpenStudy (anonymous):

Part 1: Write the general form of the equation which matches the graph below. (3 points) Part 2: In complete sentences, explain the process taken to find this equation. (3 points) The graph is below!!!

OpenStudy (anonymous):

@hba Can you help???

OpenStudy (hba):

sorry no idea

OpenStudy (anonymous):

@amistre64 @Hero Can you help???

hero (hero):

Do you know the general shape of the graph? For example, is it a parabola or a circle?

OpenStudy (anonymous):

@FoolAroundMath Help!!!!!!

hero (hero):

You don't even know the general shape of the graph?

OpenStudy (foolaroundmath):

It looks as if it's a parabola. I think the directrix (x = 4) and focus are given (-4,3). Can you find its equation?

OpenStudy (anonymous):

I'm not really good and getting the equation out of graph. Can you show me???

OpenStudy (anonymous):

@FoolAroundMath HELP!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (foolaroundmath):

The general equation of a parabola with its axis parallel to x-axis is given by \[(y-k)^{2} = 4a(x-h)\] where \((h,k)\) is the vertex of the parabola and \(a\) is the distance from the vertex to the directrix. Can you find the coordinates of the vertex? i.e. find h,k and also find the distance between the directrix and the vertex i.e. find 'a'?

OpenStudy (foolaroundmath):

edit: \[(y-k)^{2} = -4a(x-h)\] because the parabola opens up towards the left.. The formula in my prev post is valid if the parabola opens towards the right

OpenStudy (anonymous):

The vertex is (0,3).

OpenStudy (anonymous):

a = 4

OpenStudy (anonymous):

@FoolAroundMath Is it right???

OpenStudy (foolaroundmath):

Yes. So, the equation of the curve is ?

OpenStudy (anonymous):

\[(y-3)^{2} = -4(4)(x)\]

OpenStudy (anonymous):

\[(y-3)^{2} = -16x\]

OpenStudy (anonymous):

@FoolAroundMath Do you have to square (x-h)???

OpenStudy (foolaroundmath):

Nope. If the parabola had the axis to be parallel to the y-axis, the equation would have been \[(x-h)^{2} = \pm4a(y-k)\]The \(\pm\) depends on whether the parabola opens upwards or downwards (+ve for upwards)

OpenStudy (anonymous):

@FoolAroundMath So is it correct???

OpenStudy (foolaroundmath):

Yep. Your answer is correct.

OpenStudy (anonymous):

So.. PART 1: (y-3)^2 = -16x

OpenStudy (anonymous):

@FoolAroundMath So.. can you do PART 2??? Like put everything you explained to me into a small little paragraph??

OpenStudy (anonymous):

@FoolAroundMath So.. can you do PART 2??? Like put everything you explained to me into a small little paragraph??

OpenStudy (foolaroundmath):

Not my cup of coffee. Sorry. I suggest you try writing how you understood it yourself.

OpenStudy (anonymous):

@FoolAroundMath tHANX FOR ALL YOUR HELP!!! Can you help me with this question??? What is the equation of the graph? x = –1/12y2 x = 1/12y2 y = –1/12x2 y = 1/12x2

OpenStudy (foolaroundmath):

same thing, but here the equation would be \[(y-k)^{2} = 4a(x-h)\] can you find h,k,a ?

OpenStudy (anonymous):

The choices are: x = (–1/12)y^2 x = (1/12)y^2 y = (–1/12)x^2 y = (1/12)x^2